# Diffusion Generative Flow Samplers: Improving learning signals through partial trajectory optimization

### v1

@@ 5.1 Benchmarking target distributions
Funnel is a classical sampling benchmark problem from Neal (2003); Hoffman & Gelman (2011). This $10$-dimensional density is defined by $\mu(\mathbf{x})=\mathcal{N}(x^{(0)};0,9)\mathcal{N}(\mathbf{x}^{(1:9)};\mathbf{0},\exp(x^{(0)})\mathbf{I})$.
Manywell (Noé et al., 2019; Wu et al., 2020) is a $32$-dimensional distribution defined by the product of $16$ same $2$-dimensional “double well” distributions, which share the unnormalized density of $\mu(x,y)=\exp(-x^{4}+6x^{2}+0.5x-0.5y^{2})$.
@@
Baselines.
We benchmark DGFS performance against a wide range of strong baseline methods. As an MCMC methods we include the sequential Monte Carlo sampler (Moral & Doucet, 2002, SMC), which is often considered the state-of-the-art sampling method. As for normalizing flow methods we include the variational inference with normalizing flows (Rezende & Mohamed, 2015, VI-NF) approach as well as more recent de G. Matthews et al. (2022, CRAFT) which is combined with MCMC. Last but not least, we compare with PIS (Zhang & Chen, 2022) and DDS (Vargas et al., 2023a) algorithms which are closest to our framework but trained with the KL divergence formulation (Equation 5) without the improved training techniques described in Section 3.
Evaluation protocol.
@@
We display the benchmarking results in Table 1. We fail to achieve reasonable performance with DDS on Cox target where we mark it as “N/A”. The advantage of the proposed DGFS method can be observed across different tasks, indicating it could indeed benefit from its design providing more informative training signals to achieve better convergence and capturing diverse modes even in high dimensional spaces. It is also remarkable that DGFS obtains very stable results compared to other baselines. We defer related experimental details and extra experiments to Appendix C.
@@ 5.2 Analysis
Figure 5: MoG visualization of DGFS and other diffusion-based samplers shows that DGFS could capture the diverse modes well, better than PIS and DDS. The contours display the landscape of the target density.
To look into why DGFS outperforms PIS with similar architecture, we plot the learned drift network magnitude in Figure 8 in the Appendix. The result shows that the solution that DGFS achieved has more homogeneous magnitude outputs across different diffusion steps. This may explain one reason why DGFS is easier to train than PIS, since the policy probabilities at different steps $\{P_{F}(\cdot|\mathbf{x}_{n})\}_{n}$ are actually parameterized by a single neural network with temporal embedding input. This makes the learning task more stationary with DGFS. We also conduct in-depth ablation studies to evaluate the design choices and off-policy exploration ability of DGFS in Appendix C.5 and C.7.
@@ 6 Conclusion
The DGFS framework presents various opportunities for further research, fueled by its inherent limitations. To name a few, how can we better design the intermediate local signals in ways more sophisticated than the current straightforward approach? Additionally, can we harness DGFS’s off-policy exploration capabilities to succeed in high dimensional tasks, pushing the boundaries of its performance? Can we combine DGFS with a prioritized replay buffer, which is a natural fit due to its off-policy nature? This would largely reduce the number of querying $\mu(\cdot)$. From the perspective of application, there are also many interesting research questions. Can we deploy DGFS on scientific biological or chemical tasks such as protein conformation modeling? Can we incorporate equivariance into the modeling of DGFS to increase efficiency? We expect that exploring these future directions will lead to valuable insights and bring fruit follow-ups of DGFS.
@@
We provide detailed algorithmic and experimental description in Section 3 and Appendix C, and we have attached the complete code accompanying this research. To ensure robustness and reliability of our empirical results, every experiment conducted was replicated using five distinct random seeds. Upon publication, the codebase will be released on Github.
@@ C.2 $\log Z$ estimator analysis
$=\log\frac{1}{B}\sum_{b=1}^{B}\exp\left(S(\tau^{(b)})\right),\quad\tau^{(b)}\sim\mathcal{Q}.$ (41)
@@ C.4 DGFS details
DGFS contains two sets of deep network parameters: one is the same as the neural network in PIS / DDS, the other is the flow function to amortize the intermediate marginal densities. The former one is for the control network, and express it in the way of $\mathbf{f}(\cdot,n)/\sigma=\text{NN}_{1}(\cdot,n),+\text{NN}_{2}(n)\cdot\nabla\log\mu(\cdot)$. These use Fourier features to embed the step index input. We also ablate diffusion-based sampling methods without gradient information in network parametrization in C.6. The second one is a scalar output network $\text{NN}(\cdot,n)$ with similar structure. Same to PIS, we use two $64$-dimensional hidden layers after the embedding layer. For all experiments, we train with a batch size of $256$ and Adam optimizer. We have not tuned too hard on the learning rate, but simply use $1\times 10^{-4}$ and $1\times 10^{-3}$ for the policy network (i.e., drift network) and the flow function network. The training keeps for $5000$ optimization steps although almost all experiments converge in the first $1000$ steps. Similar to PIS, we set the number of diffusion steps $N$ to $100$ for all experiments. We set the diffusion step size $h$ to $0.05$ for the MoG and Cox task and $0.01$ for other tasks for all diffusion based methods. We set $\sigma$ to $1$. We set the number of evaluation particles $B$ to be $2000$ for all tasks.
Lahlou et al. (2023) also conduct toy experiments on GFlowNet-based algorithm with a control-based modeling. However, their modeling is not correct in the sense that they mistakenly use the same variance coefficient in the backward policy as in the forward policy. As shown in Equation 21, the variance coefficient at $n$-th step should be $nh\sigma^{2}/(n+1)$ rather than $h\sigma^{2}$, otherwise the forward and backward trajectory probability could not match each other. This might be one of the reasons that their simpler modeling could not scale to high dimensional problems.
@@ C.5 More experiments
We conduct ablation studies on a series of design choices of DGFS. In Figure 10, we compare the performance of DGFS and a few its variants on the VAE task. In the VAE task, the unnormalized target density $\log\mu(\cdot)$ is defined by the log latent prior $\log p(\cdot)$ plus the log decoder likelihood $\log p(\mathbf{x}_{\text{obs}}|\cdot)$, which equals the latent posterior $\log p(\cdot|\mathbf{x}_{\text{obs}})$ plus a constant value. The VAE is trained on the MNIST dataset, and $\mathbf{x}_{\text{obs}}$ is a fixed image data. We first remove the usage of the intermediate signals (i.e., the parameterization of flow function in Equation 16), which leads to slightly worse performance. We then try to train DGFS with VP modeling in Appendix B.2. For this VP modeling, we also take the form of Equation 16 for getting intermediate signals. This gives a bias of $0.86$. Notice that this is still better than both PIS and DDS, demonstrating the validity of the proposed DGFS training method. We also ablate the weighting coefficient $\lambda$ and number of diffusion steps $N$. We can see that their change would lead to small variations but the overall convergence remains good.
@@
We also compare with the FAB algorithm proposed in Midgley et al. (2022). We train FAB model with its provided open source code on the Manywell task. We try our best to tune the learning rate, batch size, $\alpha$ coefficient, number of layers and intermediate distributions. The best estimation bias result we obtain is $2.73$±$\pm 0.68$ averaged from $5$ random seeds. This is slightly worse than their reported $1.4\%$ error which corresponds to $2.306$ in log normalizing factor bias. Both the reported and our reproduced results are worse than the DGFS performance which is $0.904$ (actually, we can see that FAB is also worse than PIS and DDS). For the most challenging high dimensional Cox task, we produce a value of $28.45$±$\pm 3.55$ from the FAB model, which is also considerably worse than the performance of DGFS ($8.97$±$\pm 1.17$).
We have also tried to use a MNIST-pretrained normalizing flow as the target density. However, we find that a well-trained normalizing flow almost always assign $\infty$- ∞ to the log likelihood for out-of-distribution samples, which makes it impossible for any sampler to success. This actually makes sense because the normalizing flow is only trained on a finite size image dataset and never gets out-of-domain inputs.
@@ C.7 Off-policy exploration ability of DGFS
In Section 3.2 we have analyzed that different from other methods that learn stochastic processes for sampling, DGFS is an off-policy training algorithms. This means DGFS could train with (partial) trajectory from any distribution.
In an algorithmic perspective, we could use a slightly larger variance coefficient $\tilde{\sigma}>\sigma$ in the rollout stage of Algorithm 1 to ensure that information of more modes fall into the training data, which would drive the training of DGFS to put effort on these distant modes. To demonstrate this property, we carefully create a more difficult Mixture of Gaussian task and name it as MoG+. We intentionally put a few very distant modes in this MoG+ task, as shown by the contours in Figure 9. We set the exploration variance coefficient $\tilde{\sigma}=2$ contrary to $\sigma=1$, which achieves better mode coverage and lower normalizing factor estimation bias with the same neural network capacity. Due to limited time, we have not investigated deeper in this off-policy feature, which would be a promising future direction in the research field of probabilistic sampling.

### v3

@@ 5.1 Benchmarking target distributions
Funnel is a classical sampling benchmark problem from Neal (2003); Hoffman & Gelman (2011). This $10$-dimensional density is defined by $\mu(\mathbf{x})=\mathcal{N}(x^{(0)};0,9)\mathcal{N}(\mathbf{x}^{(1:9)};\mathbf{0},\exp(x^{(0)})\mathbf{I})$. [^7] Previous works (Zhang & Chen, 2022; Lahlou et al., 2023) unintentionally take the variance of $x^{(0)}$ to be $1$ for their methods (hence much less difficulty) but use $9$ for other methods.
Manywell (Midgley et al., 2022; Noé et al., 2019; Wu et al., 2020) is a $32$-dimensional distribution defined by the product of $16$ same $2$-dimensional “double well” distributions, which share the unnormalized density of $\mu(x,y)=\exp(-x^{4}+6x^{2}+0.5x-0.5y^{2})$.
@@
Baselines.
We benchmark DGFS performance against a wide range of strong baseline methods. As an MCMC methods we include the sequential Monte Carlo sampler (Moral & Doucet, 2002, SMC), which is often considered the state-of-the-art sampling method. As for normalizing flow methods we include the variational inference with normalizing flows (Rezende & Mohamed, 2015, VI-NF) approach as well as more recent de G. Matthews et al. (2022, CRAFT) which is combined with MCMC. We also provide results of Midgley et al. (2022, FAB) with a replay buffer that utilizes the annealed importance sampling technique to augment the training. Last but not least, we compare with PIS (Zhang & Chen, 2022) and DDS (Vargas et al., 2023a) algorithms which are closest to our framework but trained with the KL divergence formulation (Equation 5) without the improved training techniques described in Section 3.
Evaluation protocol.
We display the benchmarking results in Table 1. We fail to achieve reasonable performance with DDS on Cox target where we mark it as “N/A”. The advantage of the proposed DGFS method can be observed across different tasks, indicating it could indeed benefit from its design providing more informative training signals to achieve better convergence and capturing diverse modes even in high dimensional spaces. FAB with buffer produces highly stable results, which can be attributed to both annealed importance sampling, the use of buffers, larger networks, and other tricks. It is also remarkable that DGFS obtains very stable results compared to other baselines with diffusion modeling (PIS, DDS). We defer related experimental details and extra experiments to Appendix C.
@@ 5.2 Analysis
Figure 5: MoG visualization of DGFS and other diffusion-based samplers shows that DGFS could capture the diverse modes well. The contours display the landscape of the target density.
To look into why DGFS outperforms PIS with similar architecture, we plot the learned drift network magnitude in Figure 8 in the Appendix. The result shows that the solution that DGFS achieved has more homogeneous magnitude outputs across different diffusion steps. This may explain one reason why DGFS is easier to train than PIS, since the policy probabilities at different steps $\{P_{F}(\cdot|\mathbf{x}_{n})\}_{n}$ are actually parameterized by a single neural network with temporal embedding input. This makes the learning task more stationary with DGFS. We also conduct in-depth ablation studies to evaluate the design choices of DGFS in Appendix C.5 and C.7.
@@ 6 Conclusion
The DGFS framework presents various opportunities for further research, fueled by its inherent limitations. To name a few, how can we better design the intermediate local signals in ways more sophisticated than the current straightforward approach? Additionally, can we harness DGFS’s exploration capabilities to succeed in high dimensional tasks, pushing the boundaries of its performance? Can we combine DGFS with a prioritized replay buffer? This would largely reduce the number of querying $\mu(\cdot)$. From the perspective of application, there are also many interesting research questions. Can we deploy DGFS on scientific biological or chemical tasks such as protein conformation modeling, with equivariant modeling? We expect that exploring these future directions will lead to valuable insights and bring fruit follow-ups of DGFS.
Acknowledgement
The authors are grateful to the discussions with Francisco Vargas, Julius Berner, Lorenz Richter, Laurence Midgley, Minsu Kim, Yihang Chen, and Qinsheng Zhang.
@@ Acknowledgement
We provide detailed algorithmic and experimental description in Section 3 and Appendix C, and we have open sourced the code accompanying this research in this GitHub link.
@@ C.2 $\log Z$ estimator analysis
$\approx\log\frac{1}{B}\sum_{b=1}^{B}\exp\left(S(\tau^{(b)})\right),\quad\tau^{(b)}\sim\mathcal{Q}.$ (41)
@@ C.4 DGFS details
DGFS contains two sets of deep network parameters: one is the same as the neural network in PIS / DDS, the other is the flow function to amortize the intermediate marginal densities. The former one is for the control network, and express it in the way of $\mathbf{f}(\cdot,n)/\sigma=\text{NN}_{1}(\cdot,n),+\text{NN}_{2}(n)\cdot\nabla\log\mu(\cdot)$. These use Fourier features to embed the step index input. We also ablate diffusion-based sampling methods without gradient information in network parametrization in C.6. The second one is a scalar output network $\text{NN}(\cdot,n)$ with similar structure. Same to PIS, we use two $64$-dimensional hidden layers after the embedding layer. We set $\lambda$ to be $2$. For all experiments, we train with a batch size of $256$ and Adam optimizer. We have not tuned too hard on the learning rate, but simply use $1\times 10^{-4}$ and $1\times 10^{-3}$ for the policy network (i.e., drift network) and the flow function network. The training keeps for $5000$ optimization steps although almost all experiments converge in the first $1000$ steps. Similar to PIS, we set the number of diffusion steps $N$ to $100$ for all experiments. We set the diffusion step size $h$ to $0.05$ for the MoG and Cox task and $0.01$ for other tasks for all diffusion based methods. We set $\sigma$ to $1$. Zhang & Chen (2022) set evaluation particle number $B=6000$ for the funnel task and $B=2000$ for other tasks for unknown reasons. In this work, we set the number of evaluation particles $B$ to be $2000$ for all tasks for consistency.
@@ C.5 More experiments
We conduct ablation studies on a series of design choices of DGFS. In Figure 10, we compare the performance of DGFS and a few its variants on the VAE task. In the VAE task, the unnormalized target density $\log\mu(\cdot)$ is defined by the log latent prior $\log p(\cdot)$ plus the log decoder likelihood $\log p(\mathbf{x}_{\text{obs}}|\cdot)$, which equals the latent posterior $\log p(\cdot|\mathbf{x}_{\text{obs}})$ plus a constant value. The VAE is trained on the MNIST dataset, and $\mathbf{x}_{\text{obs}}$ is a fixed image data. We first remove the usage of the intermediate signals (i.e., the parameterization of flow function in Equation 16), which leads to slightly worse performance. We then try to train DGFS with VP modeling in Appendix B.2. For this VP modeling, we also take the form of Equation 16 for getting intermediate signals. This gives a bias of $0.86$. Notice that this is still better than both PIS and DDS, demonstrating the validity of the proposed DGFS training method. We also ablate the weighting coefficient $\lambda$ and number of diffusion steps $N$. We can see their changes lead to small variations but the overall convergence remains good.
@@
We tried to use a MNIST-pretrained normalizing flow as the target density. However, we find a well-trained normalizing flow almost always assign $\infty$- ∞ to the log likelihood for out-of-distribution samples, which makes it impossible for any sampler to success. This actually makes sense as the normalizing flow is only trained on a finite size image dataset and never gets out-of-domain inputs.
@@ C.7 Off-policy exploration ability of DGFS
The usage of the KL divergence in previous works restricts the training to be on-policy, as the training trajectory has to come from the model itself. Using off-policy samples is possible only with importance sampling, which typically results in high variance. This limits the exploration ability of these methods. On the other hand, it is first shown in Malkin et al. (2023) that GFlowNets can potentially benefit from the use of off-policy exploration. To be more specific, as has been shown by a similar analysis in prior works Malkin et al. (2023); Lahlou et al. (2023), the objectives in Equation 9 or Equation 14 do not require the training samples to follow any particular distribution (only to have full support), which means our method supports off-policy training without importance sampling. Therefore, it is possible to employ off-policy exploration by using a larger variance coefficient in the forward process during the rollout stage. In theory, this could help DGFS to better capture diversity in complex multimodal target distributions.
In an algorithmic perspective, we could use a slightly larger variance coefficient $\tilde{\sigma}>\sigma$ in the rollout stage of Algorithm 1 to ensure that information of more modes fall into the training data, which would drive the training of DGFS to put effort on these distant modes. To demonstrate this property, we carefully create a more difficult Mixture of Gaussian task and name it as MoG+. We intentionally put a few very distant modes in this MoG+ task, as shown by the contours in Figure 9. We set the exploration variance coefficient $\tilde{\sigma}=2$ contrary to $\sigma=1$, which achieves better mode coverage and lower normalizing factor estimation bias with the same neural network capacity.
We conduct an initial experiment to investigate this possibility in DGFS. We test a strategy that linearly anneals the $\tilde{\sigma}$ from $1.1$ to $1$ during the first $1000$ steps of training, and denote it as DGFS+ in Table 3. Our experiment results show that this simple strategy can improve the performance under some scenarios, and can be treated as a potential hyperparameter to extend the application scope of DGFS.
Table 3: Experimental results on benchmarking target densities. DGFS+ denotes DGFS with additional off-policy exploration trick in the initial stage of training.
MoG Funnel Manywell VAE Cox
DGFS 0.0190.019{0.019}0.019±$\pm 0.008$ 0.2740.274{0.274}0.274±$\pm 0.014$ 0.9040.904{0.904}0.904±$\pm 0.067$ 0.1800.180{0.180}0.180±$\pm 0.083$ 8.9748.974{8.974}8.974±$\pm 1.169$
DGFS+ $0.012$±$\pm 0.002$ $0.322$±$\pm 0.029$ $0.932$±$\pm 0.123$ $0.162$±$\pm 0.015$ $10.23$±$\pm 1.237$
C.8 Other ablation study
Table 4: Experimental results on benchmarking target densities for methods in Lahlou et al. (2023), with or without off-policy exploration technique. We also present performance of TB and its off-policy variant; see details in Section C.8.
MoG Funnel Manywell VAE Cox
Lahlou et al. (2023) 1.8911.891{1.891}1.891±$\pm 0.042$ 0.3980.398{0.398}0.398±$\pm 0.061$ 3.6693.669{3.669}3.669±$\pm 0.653$ 4.5634.563{4.563}4.563±$\pm 1.029$ ${2728.}$±$\pm 51.54$
Lahlou et al. (2023)+ $0.024$±$\pm 0.189$ $0.373$±$\pm 0.020$ $6.257$±$\pm 2.449$ $2.529$±$\pm 0.737$ $2722$±$\pm 31.64$
TB 0.0180.018{0.018}0.018±$\pm 0.002$ 0.3090.309{0.309}0.309±$\pm 0.023$ 1.2181.218{1.218}1.218±$\pm 0.194$ 1.0971.097{1.097}1.097±$\pm 0.239$ 138.6138.6{138.6}138.6±$\pm 7.850$
TB+ 0.0090.009{0.009}0.009±$\pm 0.003$ 0.3730.373{0.373}0.373±$\pm 0.031$ 1.2101.210{1.210}1.210±$\pm 0.192$ $\pm 0.248$ 240.4240.4{240.4}240.4±$\pm 0.998$
We run the method in Lahlou et al. (2023) in this section. As shown in Table 4, this previous method and its off-policy variant are far from comparable to the baselines in Table 1. We also run the trajectory balance method with gradient information (denoted as TB) and its variant with off-policy exploration (denoted as TB+) and present in the same table. These two methods achieve relatively better performance, but still worse than that of DGFS.